OpenStudy (anonymous):

Find the derivative of the function of y=(2x+1)/(2x-1).

OpenStudy (anonymous):

quotient rule for this one \[\left(\frac{f}{g}\right)'=\frac{gf'-fg'}{g^2}\] with \[f(x)=2x+1,f'(x)=2,g(x)=2x-1, g'(x)=2\]

OpenStudy (anonymous):

So \[\frac{ (2x-1)(2)-(2x+1)(2) }{ (2x-1)^2 }\] which simplified is \[\frac{ (4x-2)-4x-2 }{ (2x-2)^2 }\] and then \[\frac{ -4 }{ (2x-1)^2 }\]